# A Comparative Study of Competency Question Elicitation Methods from Ontology Requirements

### v1

@@ 3.2 Semantic analysis of CQ sets via sentence embeddings
To analyse the semantic characteristics of CQ sets generated by different approaches (RQ3), we conducted a study on their embeddings. This utilised Sentence-BERT embeddings from the all-MiniLM-L6-v2 model [32], which generates vectors $\mathbf{e}\in\mathbb{R}^{384}$ capturing the semantic meaning of each CQ, as done in [2, 9]].
@@
  • Coverage analysis. We measured how well one set covers the semantic content of another. This was performed in both directions, i.e. for the coverage of Set A by Set B (Set A $\leftarrow$ Set B) we compute:
    • Mean Maximum Similarity (MMS). For each CQ embedding $\mathbf{e}_{A,i}$ in Set A, its maximum cosine similarity to any CQ embedding in Set B, $s_{A_{i}\rightarrow B}=\max_{j}\cos(\mathbf{e}_{A,i},\mathbf{e}_{B,j})$, was identified. The mean of these $s_{A_{i}\rightarrow B}$ scores (and the standard deviation) indicates, on average, how well each CQ in Set A is semantically represented by its closest counterpart in Set B. A higher mean suggests stronger semantic parallels offered by Set B.
@@
  • Bidirectional coverage: This symmetric metric quantifies the overall mutual semantic overlap. It was calculated as $\frac{N_{A\rightarrow B}^{\textnormal{cov}}+N_{B\rightarrow A}^{\textnormal{cov}}}{N_{A}+N_{B}}$, where $N_{A\rightarrow B}^{\textnormal{cov}}$ is the number of CQs in Set A covered by Set B (i.e., $s_{A_{i}\rightarrow B}\geq\tau$), and $N_{B\rightarrow A}^{\textnormal{cov}}$ is the number of CQs in Set B covered by Set A. Hence, a higher percentage indicates greater shared conceptual space between the two sets.
@@ 4 Results
CQ Set # Commented Score Score $>0$ CQs
(%) ($\mu\pm\sigma$) (%)
@@
[…]
@@ 4.1 Expert evaluation results (RQ1)
CQ Set Ambiguity Relevance FKGL DCR
(%) (Score 3) ($\mu\pm\sigma$) ($\mu\pm\sigma$)
@@
Table 4: Mean and standard deviation of complexity measures per CQ set. $c_{0}$ denotes the number of characters in a CQ; $c_{1}$ quantifies the richness of ontological primitives; $c_{2}$ and $c_{3}$ measure surface-level and structural complexity, respectively (c.f. Section 3.1.2).
@@ 4.2 Computational analysis of CQ features (RQ2)
For relevance, Table 3 reports the proportion of CQs assigned a score of 3 on a 4-point Likert scale, which signifies CQs pertaining to requirements that are inferable from the user story using domain knowledge, and are functionally necessary for fulfilling the story’s goals; whilst all other CQs in this analysis achieved the maximum score of 4 (addressing explicitly stated requirements). Notably, human-annotated sets, particularly HA-2 (27.8% score 3) and HA-1 (18.2% score 3), had the highest proportions of these inferential CQs. This suggests that the ontology engineers utilised their domain expertise to elicit questions that, whilst not directly stated, capture key functional needs, such as “What is the family of an instrument?” or “What is the format of each multimedia file?” (both of which received a relevance score of 3). Conversely, Gemini had the lowest proportion of inferential (score 3) CQs (4.8%), indicating that its generated questions adhered most closely to the requirements that are explicitly stated in the user story. The Pattern (13.5%) and GPT (12.0%) sets presented an intermediate proportion of such inferential CQs.
@@
The computational readability analysis, summarised in Table 3, reveals distinct patterns across the different CQ formulation approaches. Manual CQs exhibited the highest readability; HA-1 consistently achieved the lowest scores across most indices (e.g., Flesch-Kincaid Grade Level, FKGL, $5.63\pm 2.80$), suggesting a reading level appropriate for early secondary school years. HA-2 scored slightly higher but remained highly readable (FKGL $6.88\pm 3.42$), with the lowest Dale-Chall Readability score (DCR 8.76±2.00)8.76$\pm$ 2.00)8.76 ± 2.00 ). Pattern-based CQs presented slightly reduced readability compared to the manual sets (e.g., FKGL $7.66\pm 2.81$). In contrast, LLM-generated CQ were markedly less readable, requiring comprehension levels estimated at late secondary school or early university stages. GPT yielded the lowest readability at the syllable level (FKGL $11.64\pm 2.69$); and Gemini produced slightly more readable CQs than GPT (FKGL $9.72\pm 2.67$), although the higher Dale-Chall Readability score (DCR $12.90\pm 2.64$) potentially indicates more complex or less common vocabulary usage.
@@ 5 Discussion
Our comparative analysis of CQ elicitation approaches –manual, pattern-based, and LLM-based – on AskCQ produced significant findings. The expert evaluation (RQ1) revealed a clear preference for CQs formulated by ontology engineers (HA-1, HA-2). These two sets achieved the highest suitability scores and acceptance rates (over 90%), and notably, received the least number of comments, suggesting that they had fewer perceived issues by the expert evaluators (Table 2). This preference aligns with the findings from our analysis of CQ features (RQ2). Human-generated CQs consistently demonstrated higher readability (lower Flesch-Kincaid Grade Level and Dale-Chall scores) and lower complexity across all four dimensions assessed: character length, requirement, linguistic, and syntactic complexity (Table 3 and Table 4). In contrast, LLM-generated CQs (GPT and Gemini) were generally rated as less suitable, were significantly less readable, and exhibited substantially higher complexity. Pattern-based CQs occupied a middle ground across the CQ features, but were found less suitable. Ambiguity could not be ascribed to any specific method (Table 3), and indeed, all methods generated comments that necessitated discussion among the annotators. Interestingly, the highest ambiguity was recorded for the human-annotated set HA-1, followed by Gemini and Pattern. Nevertheless, 66% of HA-1’s ambiguous CQs were resolved after a discussion phase with the annotators, as this set recorded a 91% suitability score. Instead, only 33% of ambiguous CQs from Gemini were resolved as suitable CQs after discussion. A crucial observation regarding relevance was that human CQs, particularly from HA-2, showed a higher proportion of CQs with inferential relevance (Score 3) – addressing requirements not explicitly stated in the user story but functionally necessary for its goals. This points to the role of domain knowledge or ontology engineering (OE) experience in shaping CQs.
@@
The characteristics of highly-rated human CQs identified in this study can serve as desiderata for improving automated CQ generation (e.g., via prompt engineering [37]). For instance, instructing LLMs with examples of well-formed CQs, or explicitly prompting them to consider specific properties and a desired level of semantic overlap with a controlled set of requirements, might yield outputs that are more aligned with ontology engineers. This also raises the question of what constitutes an “optimal sets” of CQs. While high overlap might indicate consensus, a degree of diversity could ensure broader coverage. Future work could explore ensemble methods for LLM-generated CQs or techniques to guide LLMs towards a more aligned, yet comprehensive, set of requirements.
Whilst the result of this study are significant, they are based on a single user story within the cultural heritage domain, and their generalisability to other domains or requirement types needs further investigation. The use of LLMs for assessing certain CQ features (relevance, components of complexity), though guided by careful prompt engineering and some manual validation, is subject to the inherent limitations of current LLM capabilities. A similar consideration applies to the complexity metrics, which, while providing comparative insights, are proxies and do not constitute a formal, exhaustive definition of CQ complexity. Hence, the value of both the LLM-based feature assessments and our complexity proxies lies in their capacity to support comparative analysis across CQ sets, rather than in the absolute range of their scores.

### v2

@@ 3.2 Semantic analysis of CQ sets via sentence embeddings
  • Coverage analysis. We measure how well one set covers the semantic content of another. This is performed in both directions, i.e. for the coverage of Set A by Set B (Set A $\leftarrow$ Set B) we compute:
    • Mean Maximum Similarity (MMS). For each CQ embedding $\mathbf{e}_{A,i}$ in Set A, its maximum cosine similarity to any CQ embedding in Set B, $s_{A_{i}\rightarrow B}=\max_{j}\cos(\mathbf{e}_{A,i},\mathbf{e}_{B,j})$, is identified. The mean of these $s_{A_{i}\rightarrow B}$ scores (and the standard deviation) indicate, on average, how well each CQ in Set A is semantically represented by its closest counterpart in Set B. A higher mean suggests stronger semantic parallels offered by Set B.
@@
  • Bidirectional coverage: This symmetric metric quantifies the overall mutual semantic overlap. It is calculated as $\frac{N_{A\rightarrow B}^{\textnormal{cov}}+N_{B\rightarrow A}^{\textnormal{cov}}}{N_{A}+N_{B}}$, where $N_{A\rightarrow B}^{\textnormal{cov}}$ is the number of CQs in Set A covered by Set B (i.e., $s_{A_{i}\rightarrow B}\geq\tau$), and $N_{B\rightarrow A}^{\textnormal{cov}}$ is the number of CQs in Set B covered by Set A. Hence, a higher percentage indicates greater shared conceptual space between the two sets.
Overall, these measures facilitate a multi-faceted comparative analysis of CQs generated from a controlled source. Although relevance metrics and general-purpose embeddings rely on parameter choices and have inherent limitations, these factors affect the results comparably across all CQ sets. Hence, we frame these computational metrics as comparative indicators of the trends characterising each generation approach rather than as absolute values.
@@ 4 Results
CQ Set # Commented (%) Score ($\mu\pm\sigma$) Score >0>0 CQs (%)
@@
In all experiments, LLMs from OpenAI and Google Cloud were prompted via their respective APIs using a configuration designed to maximise reproducibility. The parameters were fixed as follows: $\text{temperature}=0$, $\text{frequency penalty}=0$, $\text{presence penalty}=0$, and $\text{seed}=46$. This same configuration was used to generate the LLM-based CQ sets included in AskCQ (GPT4.1 and Gemini 2.5 Pro; see Section 3). All code and materials needed to reproduce our experiments are publicly available on GitHub under an MIT license.
@@ 4.1 Expert evaluation results (RQ1)
CQ Set Ambiguity (%) Relevance (% Score 3) FKGL ($\mu\pm\sigma$) DCR ($\mu\pm\sigma$)
@@
Table 4: Mean and standard deviation of complexity measures per CQ set. $c_{0}$ denotes the number of characters in a CQ; $c_{1}$ quantifies the richness of ontological primitives; $c_{2}$ and $c_{3}$ measure surface-level and structural complexity, respectively.
@@ 4.2 Computational analysis of CQ features (RQ2)
For relevance, Table 3 also reports the proportion of CQs assigned a score of 3 on a 4-point Likert scale, which signifies CQs pertaining to requirements that are inferable from the user story using domain knowledge, and are functionally necessary for fulfilling the story’s goals; whilst all other CQs in this analysis achieved the maximum score of 4 (addressing explicitly stated requirements). Notably, human-annotated sets, particularly HA-2 (27.8% score 3) and HA-1 (18.2% score 3), had the highest proportions of these inferential CQs. This suggests that the ontology engineers utilised their domain expertise to elicit questions that, whilst not directly stated, capture key functional needs, such as “What is the family of an instrument?” or “What is the format of each multimedia file?” (both of which received a relevance score of 3). Conversely, Gemini had the lowest proportion of inferential (score 3) CQs (4.8%), indicating that its generated questions adhered most closely to the requirements that are explicitly stated in the user story. The Pattern (13.5%) and GPT (12.0%) sets presented an intermediate proportion of such inferential CQs.
@@
The computational readability analysis, summarised in Table 3, reveals distinct patterns across the different CQ formulation approaches. Manual CQs exhibited the highest readability; HA-1 consistently achieved the lowest scores across most indices (e.g., Flesch-Kincaid Grade Level, FKGL, $5.63\pm 2.80$), suggesting a reading level appropriate for early secondary school years. HA-2 scored slightly higher but remained highly readable (FKGL $6.88\pm 3.42$), with the lowest Dale-Chall Readability score (DCR $8.76\pm 2.00)$. Pattern-based CQs presented slightly reduced readability compared to the manual sets (e.g., FKGL $7.66\pm 2.81$). In contrast, LLM-generated CQ were markedly less readable, requiring comprehension levels estimated at late secondary school or early university stages. GPT yielded the lowest readability at the syllable level (FKGL $11.64\pm 2.69$); and Gemini produced slightly more readable CQs than GPT (FKGL $9.72\pm 2.67$), although the higher Dale-Chall Readability score (DCR $12.90\pm 2.64$) potentially indicates more complex or less common vocabulary usage.
@@ 5 Discussion
Our comparative analysis of CQ elicitation approaches –manual, pattern-based, and LLM-based – on AskCQ produced significant findings. The expert evaluation (RQ1) revealed a clear preference for CQs formulated by ontology engineers (HA-1, HA-2). These two sets achieved the highest suitability scores and acceptance rates (over 90%), and notably, received the least number of comments, suggesting that they had fewer perceived issues by the expert evaluators (Table 2). This preference aligns with the findings from our analysis of CQ features (RQ2). Human-generated CQs consistently demonstrated higher readability (lower Flesch-Kincaid Grade Level and Dale-Chall scores) and lower complexity across all four dimensions assessed: character length, requirement, linguistic, and syntactic complexity (Tables 3 and 4). In contrast, LLM-generated CQs (GPT and Gemini) were generally rated as less suitable, were significantly less readable, and exhibited substantially higher complexity. Pattern-based CQs occupied a middle ground across the CQ features, but were found less suitable. Ambiguity could not be ascribed to any specific method (Table 3), and indeed, all methods generated comments that necessitated discussion among the annotators. Interestingly, the highest ambiguity was recorded for the human-annotated set HA-1, followed by Gemini and Pattern. Nevertheless, 66% of HA-1’s ambiguous CQs were resolved after a discussion phase with the annotators, as this set recorded a 91% suitability score. Instead, only 33% of ambiguous CQs from Gemini were resolved as suitable CQs after discussion. A crucial observation regarding relevance was that human CQs, particularly from HA-2, showed a higher proportion of CQs with inferential relevance (Score 3) – addressing requirements not explicitly stated in the user story but functionally necessary for its goals. This points to the role of domain knowledge or ontology engineering (OE) experience in shaping CQs.
@@
While the use LLMs for CQ generation is a promising research direction and a growing trend [2, 8, 31, 37], our comparative results suggest that the task requires further investigation to ensure consistent output quality. Specifically, the variability observed in our study underscores the necessity for clear instructions and carefully designed prompts regarding how to formulate CQs. In this context the characteristics of highly-rated human CQs identified in this study can serve as desiderata for improving automated CQ generation (e.g., via prompt engineering [37]). For instance, instructing LLMs with examples of well-formed CQs, or explicitly prompting them to consider specific properties and a desired level of semantic overlap with a controlled set of requirements, might yield outputs that are more aligned with ontology engineers. This also raises the question of what constitutes an “optimal sets” of CQs. While high overlap might indicate consensus, a degree of diversity could ensure broader coverage. Future work could explore ensemble methods for LLM-generated CQs or techniques to guide LLMs towards a more aligned, yet comprehensive, set of requirements.
Although the study’s results are significant, they are based on a single user story in the cultural heritage domain. Applying the approach to other domains or source materials may introduce confounding variables, making it difficult to attribute differences in CQs to the generation method rather than the source text. Therefore, its generalisability to other domains and requirement types requires further investigation.
The use of LLMs for assessing certain CQ features (relevance, components of complexity), though guided by careful prompt engineering and some manual validation, is subject to the inherent limitations of current LLM capabilities. A similar consideration applies to the complexity metrics, which, while providing comparative insights, are proxies and do not constitute a formal, exhaustive definition of CQ complexity. Hence, the value of both the LLM-based feature assessments and our complexity proxies lies in their capacity to support comparative analysis across CQ sets, rather than in the absolute range of their scores.